# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v2

@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
Here, $B(0)=+1$ is the Bernoulli number. This is similar to say that compared to the result of one plane, the effect of the (many) other planes is then to compensate for an additional $\frac{1}{2}$. This is similar to the result obtained from the geometrical analysis in Appendix B. We obtain the same result if we apply a left-handed circularly polarized light resonating with the planes with odd $j$ indices and measure the heights of the red peaks located at the two Dirac points. In that case, from the definition of $\kappa^{\prime}$ the result is $\frac{1}{2}C_{j=1}=-{\mathcal{S}}$. If the number of planes is infinite, this is equivalent as if we modify $C_{j=0}\rightarrow-C_{j=0}=C_{j=1}$ in Eq. (73). If we measure circular dichroism in this way corresponding to measure the difference of the two signals (i.e. of the left and right light polarizations) and taking the mean then this measures $\mathcal{S}$}. Now, we build a correspondence with the true measures of the light responses in the materials associated to the series ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
@@ 5 Conclusion
We have developed the formalism of local topological invariants (markers) from the magnetic monopole to lattice models with an emphasis on the Haldane model and on its topological quantum phase transition. The numerical analysis presented in this work through the various probes shows how this geometrical approach of local character in momentum space is efficient, simple and useful. The local invariants are associated to physical observables such as the quantum Hall response and local response in reciprocal space to circularly polarized light. We introduce an effective magnetic moment for the quantum monopole on the Bloch sphere such that the susceptibility response with respect to the additional magnetic field along $z$ direction is of topological character i.e. remains invariant within the topological phase. Through a proximity effect in coupled-planes systems, the QAH effect and the QSH effect are alternatively related through a even-odd effect. When taking the thermodynamical limit, we formulate a correspondence (analogy) between a $\frac{1}{2}$ invariant at a topological interface and the Ramanujan infinite alternating series. Many challenging questions remain to be addressed within this geometrical approach that we hope will be useful to the community, related to practical applications and to recent results on photoluminescence results locally resolved in momentum space [31]. Topological properties are also measurable in circuit quantum electrodynamics lattices through a local pump probe in real space [54].
@@
Here, we formulate an analogy with a charge $e$ moving from north to south in an electric field oriented along the unit vector related to the dressed polar angle. The second Newton equation is $\hbar\dot{\tilde{\theta}}=e{\mathcal{E}}$. From quantum mechanics and Parseval-Plancherel theorem, it is then possible to show the existence of a transverse pumped current of the form [4, 24]
@@
  • [49] K. Le Hur, S. A. Saati, “Quantum Hall and Light Responses in a 2D Topological Semimetal”, Comptes Rendus de Physique, Académie des Sciences 25 (2024), p. 415-432.

### v3

@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
Here, $B(0)=+1$ is the Bernoulli number. This is similar to say that compared to the result of one plane, the effect of the (many) other planes is then to compensate for an additional $\frac{1}{2}$. This is similar to the result obtained from the geometrical analysis in Appendix B. We obtain the same result if we apply a left-handed circularly polarized light resonating with the planes with odd $j$ indices and measure the heights of the red peaks located at the two Dirac points. In that case, from the definition of $\kappa^{\prime}$ the result is $\frac{1}{2}C_{j=1}=-{\mathcal{S}}$. If the number of planes is infinite, this is equivalent as if we modify $C_{j=0}\rightarrow-C_{j=0}=C_{j=1}$ in Eq. (92). If we measure circular dichroism in this way corresponding to measure the half difference of the two signals (i.e. of the left and right light polarizations) then this measures $\mathcal{S}$}. Now, we build a correspondence with the measures of the light responses in the materials associated to the series ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
@@ 5 Conclusion
We have developed the formalism of local topological invariants (markers) from the magnetic monopole to lattice models with an emphasis on the Haldane model and on its topological quantum phase transition. The numerical analysis presented in this work through the various probes shows how this geometrical approach of local character in momentum space is efficient, simple and useful. The local invariants are associated to physical observables such as the quantum Hall response and local response in reciprocal space to circularly polarized light. We introduce an effective magnetic moment for the monopole such that the susceptibility response with respect to the additional magnetic field along $z$ direction reveals the topological phase transition. Through a proximity effect in coupled-planes systems, the QAH effect and the QSH effect are alternatively related through a even-odd effect. When taking the thermodynamical limit, we formulate a correspondence (analogy) between a $\frac{1}{2}$ invariant at a topological interface and the Ramanujan infinite alternating series. Many challenging questions remain to be addressed within this geometrical approach that we hope will be useful to the community, related to practical applications and to recent results on photoluminescence results locally resolved in momentum space [31]. Topological properties are also measurable in circuit quantum electrodynamics lattices through a local pump probe in real space [56].
@@
Here, we formulate an analogy with a charge $e$ navigating from north to south in an electric field oriented along the unit vector related to the dressed polar angle. The second Newton equation is $\hbar\dot{\tilde{\theta}}=e{\mathcal{E}}$. From quantum mechanics and Parseval-Plancherel theorem, it is then possible to show the existence of a transverse pumped current of the form [4, 24]
@@
  • [47] M. Mogi, Y. Okamura, M. Kawamura, R. Yoshimi, K. Yasuda, A. Tsukazaki, K. Takahashi, T. Morimoto, N. Nagaosa, M. Kawasaki, “Experimental signature of the parity anomaly in a semi-magnetic topological insulator”, Nat. Phys. 18 (2022), p. 390.
  • [48] D. Zhuo, B. Zhang, H. Zhou, H. Tay, X. Liu, Z. Xi, C.-Z. Chen, C.-Z. Chang, “Half-Quantized Chiral Edge Current in a C=1/2 Parity Anomaly State”, Phys. Rev. Lett. 136 (2026), p. 016601.
@@
  • [54] G. Montambaux, “Ramanujan, Landau and Casimir, divergent series : a physicist approach”, Comptes Rendus. Physique, Académie des Sciences 26 (2025), p. 515-531.